OpenStudy (anonymous):

when the cost of a car is multiplied by 0.08 the result is 1770?

OpenStudy (anonymous):

\[Cost * 0.08 = 1770\] can you find cost now?

hartnn (hartnn):

Hi, @Wisselly \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Do you need to find cost of car ? if so, then C*0.08=1770 just divide by 0.08 on both sides, and you'll get C = cost of the car.

hartnn (hartnn):

Did you try ? what did you get the cost as ??

OpenStudy (anonymous):

221.24

hartnn (hartnn):

That must be 1770/8, right ? so, Cost=1770/0.08 = 22125 . got it ?